A uniform meter stick is supported on a knife edge
Answer #1

knife edge at the 32 cm mark

. when a 120 g mass is hung at the 16 cm?

1m = 100cm

Let: R = reaction force @ the knife support
W = weight of the meter stick

Sum of moments @ 0cm mark is equal to 0, assume clockwise rotation is positive:

(120 g)(16 cm) - R(32 cm) + W(50 cm) = 0

120 * (32 -16) = W * (50 - 32)

-> W = 120 * 16 / 20 = 96 gram